OpenStudy (maeganr):

50 + 25x = y 35 + 27.50X = y Does anyone know how to solve this? I am clueless.

OpenStudy (anonymous):

A system of equations, you see you have y solved for in the first equaton, so you can take that and plug it in for y in the other equation...then theres only x to solve for so y=50+25x plug 50+25x into the other equation where y is 35+27.50x=50+25x now solve for x

OpenStudy (maeganr):

Is it... 27.50x = 15 + 25x 22.50x = 15 x= 22.5/15

OpenStudy (maeganr):

I combined to get the x on one side and the whole numbers on the other side

OpenStudy (anonymous):

you should get 6... 27.50x=15+25x 2.5x=15 x=15/2.5 x=6

OpenStudy (anonymous):

Now that you know x is 6 plug that into one of the original equations and solve for y

OpenStudy (maeganr):

ok

OpenStudy (maeganr):

50 + 25(6) = y 50+ 150 = y Y = 200 is this right?

OpenStudy (anonymous):

Yep

OpenStudy (maeganr):

Yeah!

OpenStudy (maeganr):

Thank you.

OpenStudy (anonymous):

You're welcome